# Dependency-Aware Semi-Structured Sparsity of GLU Variants in Large Language Models

### v1

@@ 4.1 Settings
DaSS’s performance is evaluated over popular LLMs using SwiGLU, including the LLaMA2 model family (Touvron et al., 2023b), which has models with parameters ranging between 7 billion and 70 billion, and also the Mistral-7B model (Jiang et al., 2023). Among them, LLaMA2-70B and Mistral-7B use grouped-query attention (Ainslie et al., 2023), the MLP module parameter accounts for around 80% of the total model parameters. To test the generalization ability to ReGLU models, we use ReluLLaMA (Team, 2023), which is fine-tuned using ReGLU variant (Shazeer, 2020; Mirzadeh et al., 2023) based on LLaMA2 with small accuracy loss. The model configuration details are in Appendix 7. We access the public checkpoints of the involved models provided by the HuggingFace Transformers library (Wolf et al., 2019).
@@
As shown in Sun et al. (2023), the pre-LLM methods become less effective for LLMs. We compare the performance with two LLM-specific one-shot pruning approaches, SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023). Those baseline methods utilize uniform layerwise sparsity that can be easily converted into hardware-friendly N:M sparsity pattern. We used the same calibration data set as SparseGPT and Wanda in their model pruning processes, consisting of 128 sequences of 2048 tokens each, randomly selected from the first shard of the C4 dataset (Raffel et al., 2020).
@@
To comprehensively evaluate the efficacy of our proposed method, three unique metrics are utilized to evaluate the performance of the pruned models: (1) perplexity (PPL) of language modeling (2) zero-shot accuracy on 5 commonsense reasoning tasks (3) 5-shot accuracy on Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021). Although perplexity has been regarded as a consistent and reliable metric for measuring compressed models in previous works (Dettmers & Zettlemoyer, 2023; Frantar & Alistarh, 2023), a recent study indicates that perplexity fails to capture the change in capabilities of compressed LLMs on performing knowledge-intensive MMLU tasks (Jaiswal et al., 2023). For perplexity evaluation, we use the validation dataset of WikiText2 (Merity et al., 2017). For zero-shot commonsense reasoning tasks, we choose five widely used tasks for accuracy evaluation: ARC (Easy and Challenge) (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PiQA (Bisk et al., 2020), and WinoGrande (Sakaguchi et al., 2021), implemented in the Lm-Evaluation-Harness (Gao et al., 2021). For 5-shot MMLU evaluation, we use Chain-of-Thought Hub (Fu et al., 2023) which is based on the official implementation of MMLU (Hendrycks et al., 2021). The MMLU encompasses 57 tasks, spanning from STEM, Humanities, Social Sciences, among others and we report the mean accuracy of 57 tasks. We evaluate the (1) perplexity of all the aforementioned models. To fully demonstrate the task-wise performance in different sparsity patterns, we report the (2) and (3) task performance of the largest LLaMA2-70B model.
Table 2: WikiText perplexity of pruned LLaMA-2 and Mistral models
Model Family Mistral Llama2
Size 7B 7B 13B 70B
Method MLP Sparsity PPL ($\downarrow$)
Dense - 5.25 5.47 4.88 3.32
SparseGPT 4:8 7.36 7.33 6.29 4.66
Wanda 4:8 7.38 7.63 6.42 4.59
DaSS 4:8 7.06 7.26 6.16 4.41
SparseGPT 2:4 8.86 8.72 7.30 5.32
Wanda 2:4 9.24 9.55 7.68 5.35
DaSS 2:4 8.39 8.48 6.90 4.91
SparseGPT 50% 6.20 6.38 5.60 4.06
Wanda 50% 6.25 6.50 5.67 4.07
DaSS 50% 6.15 6.44 5.59 4.00
Table 3: Downstream tasks performance of LLaMA2-70B model in different sparsity pattern. The MMLU scores of dense LLaMA2-70B and LLaMA2-34B reported in Touvron et al. (2023b) are 68.9 and 62.6 respectively.
Sparsity Methods Commonsense Reasoning (0 shot) ($\uparrow$) World Knowledge ($\uparrow$)
PIQA HellaSwag Winogrande ARC-e ARC-c Average MMLU (5 shot)
Dense - 82.15 66.05 77.98 82.55 54.35 72.62 69.10
4:8 SparseGPT 80.52 61.00 77.03 79.85 50.60 69.80 60.24
Wanda 80.47 61.85 75.45 80.10 50.00 69.57 59.69
DaSS 80.79 62.70 76.09 81.20 51.19 70.39 60.86
DaSS+skip 1/4 81.34 63.10 76.09 80.05 50.68 70.25 65.82
2:4 SparseGPT 79.00 59.00 76.64 78.50 47.87 68.20 56.99
Wanda 79.22 59.25 74.66 78.90 47.01 67.81 56.41
DaSS 79.70 60.00 74.82 79.65 49.15 68.66 57.53
DaSS+skip 1/4 80.63 62.05 74.51 80.00 49.40 69.32 64.36
50% SparseGPT 81.50 64.20 78.45 81.90 52.73 71.76 64.52
Wanda 81.01 64.30 77.35 80.95 52.05 71.13 62.72
DaSS 81.18 64.60 77.90 81.35 51.71 71.35 63.27
DaSS+skip 1/4 81.23 64.45 77.27 81.35 52.81 71.42 66.81
@@
In the less interpretable perplexity evaluation, we only prune the MLP layers. In the task-wise evaluation of LLaMA2-70B model, we prune both attention and MLP modules in accordance with previous works to better understand the performance gap between the pruned and original models. For attention module pruning, we use Wanda method, which is more efficient than SaprseGPT. In LlaMA2-70B, as the MLP module accounts for more than 80% total parameters, we observe the choice of attention pruning has almost no impact on the final performance. We apply a uniform sparsity ratio across all the pruned layers and conduct evaluations using three sparsity types: unstructured sparsity, and semi-structured sparsities of 4:8 and 2:4.
@@ 4.2 Language Modeling
This evaluation involved a detailed examination of all the opened LLaMA2 models and Mistral 7B model. In perplexity evaluation, as shown in Table 2, our method consistently achieves better performance than SparseGPT and Wanda in more constrained and practical N:M sparsity pattern. As indicated by Sun et al. (2023), where weight updates can improve the performance in N:M sparsity pattern, our method shows superior performance even without computationally expensive weight updates. For Mistral 7B and LLaMA2 70B models with larger MLP layers, our method also outperforms SparseGPT in unstructured sparsity.
@@ 4.3 Downstream Tasks
While perplexity has been a popular metric in earlier LLM compression works (Frantar & Alistarh, 2023; Dettmers & Zettlemoyer, 2023), it essentially measures the confidence of a language model in text prediction and doesn’t always reflect its proficiency in performing downstream tasks (Jaiswal et al., 2023). Apart from assessing perplexity, we comprehensively evaluate the performance of pruned LLaMA2-70B models in downstream tasks.
Comparison with Baselines. In Table 3, we present the performance of different sparse LLaMA2-70B models on downstream tasks with prompting. The results show that our method outperforms SparseGPT and Wanda in most tasks at semi-structured N:M sparsity pattern including the aggregated task MMLU. The only exception is that SparseGPT outperforms both Wanda and DaSS in Winogrande task. For unstructured sparsity, our method outperforms Wanda sharing the same complexity level. It is noteworthy that the improvement of DaSS over Wanda becomes more pronounced in the challenging MMLU task, where it achieves an increase in accuracy of 1.17 and 1.12 for the 4:8 and 2:4 sparsity patterns, respectively.
Partial N:M Sparsity & Layer Sensitivity. In accordance with the observations with Jaiswal et al. (2023), we see a considerable performance degradation in knowledge-intensive MMLU task, in which only unstructured 50% sparsity models outperform the dense LLaMA2-34B model. As GPU plays a more important role in larger model inference, it is essential to improve the performance of pruned models in hardware-friendly N:M sparsity. Uniform reduction of the overall sparsity level is not feasible for N:M sparsity, Frantar & Alistarh (2023) suggests a specific subset of layers can be chosen for full N:M sparsification. Here we skip pruning 1/4 consecutive layers (20 layers) and result in a final 37.5% sparsity ratio. We use the first 10 tasks of MMLU to study pruning sensitivity. We divide the model into 4 consecutive parts and study skipping pruning each part. As shown in Table 4, the earlier layers are more sensitive than the later ones in knowledge-intensitive tasks, which is contradictory to the findings in Frantar & Alistarh (2023) using perplexity.
Start Index 0 10 20 40 60 Dense
Acc (%) 60.12 61.72 59.75 56.54 56.83 64.86
An extensive line of work (Haviv et al., 2023; Meng et al., 2022; Geva et al., 2023) discussed how factual knowledge is stored in Transformer parameters, and early to middle MLP sublayers are demonstrated to be crucial for memorized predictions. Interestingly, the layer sensitivity results have a strong correlation with the findings in those works. We also apply the searched layer index in MMLU tasks to commonsense reasoning tasks, and we see fewer performance gains, which highlights the importance of task-specific pruning.
4.4 Generalization to ReGLU
We test the generalization ability to another GLU variant, ReGLU. In Table 5, we also observed DaSS outperform SparseGPT and Wanda in N:M sparsity pattern. This demonstrates the generalization ability of our method to other GLU variants. Since augmenting input activations is important for the expanding projection of non-GLU MLP modules, our methods cannot achieve better performance than Wanda in older LLMs that do not use GLU variants.
Table 5: WikiText perplexity of ReluLLaMA models
Size 7B 13B
Method MLP Sparsity PPL ($\downarrow$)
Dense - 6.15 5.50
SparseGPT 4:8 8.44 7.45
Wanda 4:8 9.21 7.74
DaSS 4:8 8.19 7.16
SparseGPT 2:4 10.26 8.98
Wanda 2:4 12.68 9.87
DaSS 2:4 9.61 8.18
SparseGPT 50% 7.22 6.39
Wanda 50% 7.52 6.53
DaSS 50% 7.24 6.40
@@ 4.5 Performance Analysis
Sparsity Variation.
Figure 2 illustrates a comparison of the mean zero-shot task accuracy at varying levels of sparsity for the MLP component of the LLaMA2-70B model. It’s evident that DaSS pruning maintains competitive performance which closely matches that of SparseGPT across the entire range of sparsity ratios tested. Notably, DaSS achieves such performance without the need for weight updates, suggesting its effectiveness and efficiency in locating sparse neural networks. On the other hand, output-balanced Wanda pruning shows a significant decline in accuracy as the sparsity ratio increases. This suggests that Wanda pruning may suffer from structural mismatch issues within the MLP layer, which become more pronounced at higher sparsity levels. As a result, the neural network’s performance deteriorates, potentially leading to a dysfunctional model at extreme sparsity ratios.
Figure 2: Mean zero-shot tasks performance at different MLP sparsity ratios.
Robustness to calibration samples.
Figure 3: Robustness to calibration samples.
@@ 4.6 Running Time Analysis
As observed in our preceding study on layer sensitivity, the efficiency of the pruning process is crucial when numerous iterations are necessary to reach the desired performance for specific tasks. For DaSS and Wanda, the computational complexity is quantified as $O(d^{2})$, while SparseGPT exhibits a higher complexity of $O(d^{3})$. We recorded the overall time taken for pruning MLP layers, not including the forward pass process, in accordance with the approach described by Sun et al. (2023). We use a single A6000 48GB GPU to prune the 7B and 13B models, and use 8 A100 40GB GPUs to prune the larger 70B model.
As demonstrated in Table 6, the computational overhead incurred by DaSS is minimal, especially when compared to SparseGPT. While the processing speed of DaSS is marginally slower than that of Wanda, this can be attributed to Wanda’s more efficient approach of sorting weights exclusively along the last dimension. Nonetheless, given the substantial improvements in accuracy achieved by DaSS, this additional computational time is justifiable and beneficial.
Table 6: Pruning speed (in seconds) comparison.
@@
5 Related Work
Pruning LLM. Neural network pruning in LLM can be broadly categorized into two groups: structured pruning (Ma et al., 2023; Zhang et al., 2023) and unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023). Ma et al. (2023) proposes a dependency detection algorithm to detect and prune non-critical grouped structures followed by LoRA fine-tuning. Although structured pruning can usually have better hardware efficiency, the accuracy drops a lot even at a low compression rate. Unstructured pruning can yield a higher compression rate and achieve acceleration on Nvidia’s GPUs by employing a hardware-friendly N:M sparsity pattern. SparseGPT (Frantar & Alistarh, 2023) leverages the Hessian inverse for pruning and reduces reconstruction error of dense and sparse weights by subsequent weight updates. Wanda (Sun et al., 2023) employs an efficient method that augments input activations into weight magnitudes, and matches the performance of SparseGPT at medium sparsity. Our work incorporates dependency information into unstructured pruning, achieving a novel pruning paradigm.
Inherent Sparsity of Transformer MLP. Interestingly, sparsity within the MLP activations of trained Transformer-based models occurs innately even without applying explicit regularizations or constraints (Zhang et al., 2022; Li et al., 2023; Dong et al., 2023). Such a phenomenon is prevalent in learned Transformers, including other zero-saturating functions. Liu et al. (2023); Mirzadeh et al. (2023); Zhang et al. (2022) achieve actual LLM inference speedup by only performing computation corresponding to the activating neuron for a given input. They do not actually reduce the model size since they mainly reduce I/O and computation latency in a selective weights loading manner, and thus, these methods are less applicable in large batch-size inference settings. Our work investigates the weight sparsity in MLP module by considering corresponding intermediate activations.
Outlier-dependent LLM Compression. Outlier features, defined as features with magnitudes substantially larger than others, are a notable characteristic of LLMs (Dettmers et al., 2022). Despite making up only a small fraction of all feature dimensions, these outliers play a critical role in attention and predictive performance. Such observation has impeded the development of some LLM-specific quantization methods (Dettmers et al., 2022; Xiao et al., 2023; Lin et al., 2023; Ashkboos et al., 2023) to handle outliers more effectively. Wanda (Sun et al., 2023) broadens these insights, revealing that outlier features are significant in deciding weight importance when pruning LLMs. Our analysis diverges from conventional wisdom by demonstrating that, in the context of SwiGLU-based MLP input projections, the influence of outlier features is not as pronounced as previously assumed, prompting a reevaluation of their role in LLM compression strategies.
@@ 6 Conclusion
We propose the Dependency-aware Semi-structured Sparsity (DaSS) method which effectively addresses the challenges of pruning SwiGLU-based MLP modules LLMs. DaSS strikes a unique balance between the adaptability of unstructured pruning and the orderliness of structured pruning. By leveraging the MLP intermediate activation norms as the group importance indicator, we develop a novel pruning metric that assesses weight importance in a more structurally consistent manner. Empirical evaluations on the Mistral and LLaMA2 model families demonstrate that DaSS surpasses state-of-the-art LLM pruning methods like SparseGPT and Wanda in achieving hardware-friendly N:M sparsity patterns. Our research further reveals that selectively skipping the pruning of earlier layers can lead to significant performance improvements in knowledge-intensive tasks. The efficiency of the pruning process becomes paramount when multiple iterations are essential to tailor and optimize the sparsity of LLMs for varied tasks. This underscores the potential of our DaSS method in facilitating rapid and task-specific model pruning to achieve optimal performance.
Impact Statements
The present study is focused on advancing the field of large language model deployment with an emphasis on efficiency. While there are numerous societal implications of our work, we believe none require specific emphasis in this context.
Appendix A Model Configurations

### v2

@@ Discussion.
PPL ($\downarrow$)
MLP Sparsity Method Gemma (GeGLU) ReluLLaMA (ReGLU) Mistral (SwiGLU) LLaMA2 (SwiGLU)
7B 7B 13B 7B 7B 13B 70B
Dense - 7.00 6.15 5.50 5.25 5.47 4.88 3.32
4:8 SparseGPT 10.86 8.44 7.45 7.36 7.33 6.29 4.66
Wanda 14.41 9.21 7.74 7.38 7.63 6.42 4.59
DaSS 10.86 8.19 7.16 7.06 7.26 6.16 4.41
2:4 SparseGPT 13.93 10.26 8.98 8.86 8.72 7.30 5.32
Wanda 32.50 12.68 9.87 9.24 9.55 7.68 5.35
DaSS 13.69 9.61 8.18 8.39 8.48 6.90 4.91
50% SparseGPT 9.64 7.22 6.39 6.20 6.38 5.60 4.06
Wanda 9.93 7.52 6.53 6.25 6.50 5.67 4.07
DaSS 9.11 7.24 6.40 6.15 6.44 5.59 4.00
Table 2: WikiText perplexity of pruned LLMs. The (*) in Size indicates a model that uses larger MLP than Vallina Transformer (e.g., grouped query attention).
@@ 5.1 Settings
DaSS’s performance is evaluated over open LLMs using GLU variants. SwiGLU is the most widely used GLU-based MLP in the recent LLMs, including the LLaMA2 model family Touvron et al. (2023), which has models with parameters ranging between 7 billion and 70 billion, and also the Mistral-7B model Jiang et al. (2023). Among them, LLaMA2-70B and Mistral-7B use grouped-query attention Ainslie et al. (2023), the MLP module parameter accounts for around 80% of the total model parameters. There are only a few open LLMs that use other variants. For GeGLU, we use Gemma-7B. It is worth noting that the MLP intermediate dimension of Gemma-7B is 8×8$\times$8 × model dimension, making the MLP module much larger. For ReGLU, we use ReluLLaMA Team (2023), which is fine-tuned using ReGLU variant Shazeer (2020); Mirzadeh et al. (2023) based on LLaMA2 with small accuracy loss. The model configuration details are in Appendix A.1. We access the public checkpoints of the involved models provided by HuggingFace Transformers Wolf et al. (2019).
@@
We compare the performance with two LLM-specific one-shot pruning approaches, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024). We don’t consider structured pruning methods like LLM-Pruner Ma et al. (2023) and Sheared LLaMA Xia et al. (2023b), as they typically require re-training to recover accuracy, and are less practical for large models like LLaMA2-70B. Those baseline methods utilize uniform layerwise sparsity that can be easily converted into hardware-friendly N:M sparsity pattern. We used the same calibration data set as SparseGPT and Wanda in their model pruning processes, consisting of 128 sequences of 2048 tokens each, randomly selected from the first shard of the C4 dataset Raffel et al. (2020).
@@
To comprehensively evaluate the efficacy of our proposed method, two unique metrics are utilized to evaluate the performance of the pruned models: (1) perplexity (PPL) of language modeling (2) zero-shot accuracy on 5 commonsense reasoning tasks. Perplexity has been regarded as a consistent and reliable metric for measuring compressed models Dettmers and Zettlemoyer (2023); Frantar and Alistarh (2023), downstream tasks tend to be noisy, but more interpretable. For perplexity evaluation, we use the validation dataset of WikiText2 Merity et al. (2017). For zero-shot commonsense reasoning tasks, we choose five widely used tasks for accuracy evaluation: ARC (Easy and Challenge) Clark et al. (2018), HellaSwag Zellers et al. (2019), PiQA Bisk et al. (2020), and WinoGrande Sakaguchi et al. (2021), implemented in the Lm-Evaluation-Harness Gao et al. (2021). We evaluate the perplexity of all the aforementioned models. To fully demonstrate the task-wise performance in different sparsity patterns, we report the downstream task performance of the largest LLaMA2-70B model.
@@
In the less interpretable perplexity evaluation, we only prune the MLP layers. In the task-wise evaluation of LLaMA2-70B model, we prune both attention and MLP modules in accordance with previous works to better understand the performance gap between the pruned and original models. For attention module pruning, we use Wanda method, which is more efficient than SparseGPT. In LlaMA2-70B, as the MLP module accounts for more than 80% total parameters, we observe the choice of attention pruning has almost no impact on the final performance. We apply a uniform sparsity ratio across all the pruned layers and evaluate three sparsity types: unstructured sparsity, and semi-structured sparsities of 4:8 and 2:4.
@@ 5.2 Language Modeling
This evaluation involved a detailed examination of all the aforementioned LLMs. In perplexity evaluation, as shown in Table 2, our method consistently achieves better performance than SparseGPT and Wanda in more constrained and practical N:M sparsity pattern. As indicated by Sun et al. (2024), where weight updates can improve the performance in N:M sparsity pattern, our method shows superior performance even without computationally expensive weight updates. For Mistral-7B, Gamma-7B and LLaMA2-70B models with larger MLP layers, our method also outperforms SparseGPT in unstructured sparsity. Impressively, for the Gemma-7B with much larger MLP layers, DaSS outperforms Wanda in N:M sparsity significantly. DaSS shows consistent effectiveness across SwiGLU, GeGLU, and ReGLU, proving the generalizability of the DaSS method on GLU variants.
@@ 5.3 Downstream Tasks
While perplexity has been a popular metric in earlier LLM compression works Frantar and Alistarh (2023); Dettmers and Zettlemoyer (2023), its result is less interpretable. Apart from assessing perplexity, we comprehensively evaluate the performance of pruned LLaMA2-70B models in the widely used commonsense reasoning tasks.
In Table 3, we present the performance of different sparse LLaMA2-70B models on downstream tasks with prompting. The results show that our method outperforms SparseGPT and Wanda in most tasks at semi-structured N:M sparsity pattern. The only exception is that SparseGPT outperforms both Wanda and DaSS in Winogrande task. 50% unstructured SparseGPT pruned model even outperforms the dense model. Such results align with the prevailing LLM compression works Dettmers and Zettlemoyer (2023); Frantar and Alistarh (2023), where single task results can be noisy, the mean accuracy provides stable and reliable results. For unstructured sparsity, our method outperforms Wanda sharing the same complexity level.
Sparsity Method CommonSenseQA ($\uparrow$)
PIQA HellaSwag Winogrande ARC-e ARC-c Average
Dense - 82.15 66.05 77.98 82.55 54.35 72.62
4:8 SparseGPT 80.52 61.00 77.03 79.85 50.60 69.80
Wanda 80.47 61.85 75.45 80.10 50.00 69.57
DaSS 80.79 62.70 76.09 81.20 51.19 70.39
2:4 SparseGPT 79.00 59.00 76.64 78.50 47.87 68.20
Wanda 79.22 59.25 74.66 78.90 47.01 67.81
DaSS 79.70 60.00 74.82 79.65 49.15 68.66
50% SparseGPT 81.50 64.20 78.45 81.90 52.73 71.76
Wanda 81.01 64.30 77.35 80.95 52.05 71.13
DaSS 81.18 64.60 77.90 81.35 51.71 71.35
Table 3: Downstream tasks performance of LLaMA2-70B model in different sparsity pattern.
@@ 5.4 Performance Analysis
Sparsity Variation
MLP Sparsity 40% 50% 60% 70% 80%
SparseGPT 72.28 71.97 70.08 65.53 52.14
Wanda 72.35 71.49 69.47 62.75 40.76
DaSS 72.51 71.89 70.05 66.38 53.00
Table 4: LLaMA2-70B mean zero-shot tasks performance at different MLP sparsity ratios
Table 4 illustrates a comparison of the mean zero-shot task accuracy at varying levels of sparsity for the MLP module of the LLaMA2-70B model. It’s evident that DaSS pruning maintains competitive performance which closely matches that of SparseGPT across the entire range of sparsity ratios tested. Notably, DaSS achieves such performance without the need for weight updates, suggesting its effectiveness and efficiency in locating sparse neural networks. On the other hand, outlier-dependent Wanda pruning method shows a significant decline in accuracy as the sparsity ratio increases. This clear empirical evidence challenges the prevailing notion Dettmers et al. (2022); Lin et al. (2024) that weights corresponding to outliers are more important.
@@ 5.5 Running Time Analysis
For DaSS and Wanda, the computational complexity is quantified as $O(d^{2})$, while SparseGPT exhibits a higher complexity of $O(d^{3})$. We recorded the overall time taken for pruning MLP layers, not including the forward pass process, in accordance with the approach described by Sun et al. (2024). We use a single A6000 48GB GPU to prune the 7B and 13B models, and use 8 A100 40GB GPUs to prune the larger 70B model.
@@
Table 5: Pruning speed (in seconds) comparison.
As demonstrated in Table 5, the computational overhead incurred by DaSS is minimal, especially when compared to SparseGPT. Although the processing speed of DaSS is marginally slower than Wanda, this difference arises primarily from the implementation details of torch.sort(). In the implementation of torch.sort(), sorting along the last dimension is more efficient. Therefore, the observed delay in DaSS is attributed to the sorting operations rather than an inherent inefficiency in our method. Despite this, the substantial improvements in accuracy achieved by DaSS make the additional computational time justifiable and beneficial.
Sparsity Dense 50% 70%
Tokens per seconds 3.05 5.64 8.27
Speedup 1.0×1.0$\times$1.0 × 1.85×1.85$\times$1.85 × 2.71×2.71$\times$2.71 ×
Table 6: Decode inference performance for varying sparsity levels with LLaMA2-7B utilizing DeepSparse NeuralMagic (2021) inference engine.
Inference efficiency
We prune both attention and MLP module to test speed up. The speedup achieved by sparse model, as demonstrated in Table 6, highlights a significant reduction in end-to-end decode latency when utilizing LLaMA2-7B on the DeepSparse inference engine NeuralMagic (2021). The speed is tested on an Intel Xeon Platinum 8160 CPU with 24 cores. We highlight that the inference speedup is not exclusive to our pruning method but is a result of the inherent power of sparsity in accelerating computation.
@@ 6 Conclusion
Our observations reveal a diminishing importance of outliers in GLU-based MLP modules. Based on this insight, we propose the Dependency-aware Semi-structured Sparsity (DaSS) method. DaSS effectively balances the adaptability of unstructured pruning with the structural consistency of dependency-based structured pruning, utilizing MLP intermediate activation norms as the group importance indicator to develop a novel pruning metric. Empirical evaluations on the Mistral, Gemma and LLaMA2 model families show that DaSS surpasses state-of-the-art pruning methods such as SparseGPT and Wanda in achieving hardware-friendly N:M sparsity patterns. We hope our work inspires further research and development in the compression of GLU-based LLMs.
Limitations
Our research has the following limitations:
Merits beyond inference efficiency
Beyond improving inference efficiency, pruning techniques have demonstrated additional benefits for LLMs. Specifically, pruned models tend to exhibit fewer hallucinations compared to their unpruned counterparts Chrysostomou et al. (2024). Additionally, pruning has been shown to enhance resilience against jailbreak prompting Hasan et al. (2024) and to help in reducing toxicity and bias Xu and Hu (2022). However, the potential advantages of DaSS pruning beyond inference efficiency remain unclear and warrant further investigation.
Speed test in GPU
For LLM inference, the primary contributor to inference latency is the token generation process. While PyTorch’s current 2:4 sparsity implementation is optimized for accelerating matrix multiplication, it does not significantly enhance the performance of token generation. Recent work nm-vllm Shaw and Goin (2024) has shown approximately 1.5×1.5$\times$1.5 × token generation speedup for LlaMA2-7B at 50% unstructured sparsity. However, since this library is newly released and still under active development, it currently has many bugs, and we have not yet successfully implemented it. Kurtic et al. (2023) reports the speedup of several N:M sparsity patterns, its implementation hasn’t been released yet. The hardware and software support for sparsity is rapidly advancing, which will make sparse models more user-friendly in the future.
Application to other model architectures
Our method leverages a newly discovered property of GLU-based MLPs, making it unsuitable for older LLMs, such as OPT Zhang et al. (2022), which do not incorporate GLU variants. Additionally, our approach does not extend to the attention module, as the outlier activations still play an important role in the attention module.
Appendix A Appendix
A.1 Model Configurations